# On the strong DR/DZ equivalence conjecture

### v1

@@ 4.2. Step $1$: the key combinatorial lemma
We describe a result in [4] in quite general terms adapted to the needs of our argument.
@@
As we see from the statement of Lemma 4.1, is it more convenient to use the class $\tilde{\Xi}^{m}_{g,n}\coloneqq\sum_{d\geq 0}(-1^{d})(\Xi^{m}_{g,n})_{d}$.
@@ 4.3. Step $2$: equivalence between the master relation and DR/DZ
It follows by induction using the key lemma that the master relation is equivalent to Conjecture 2.1, more precisely:
@@ 4.4. Step 3: proof of the master relation in the Gorenstein quotient
We prove Theorem 3.5 for $\alpha=\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}$ which is implies Theorem 3.5 for any $\alpha$ by Lemma 3.6.
@@
    (4.4) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+1}}\psi_{n+1}^{b}\left(B^{1}_{g^{\prime},n^{\prime}}-A^{1}_{g^{\prime},n^{\prime}}\right)\leq 2g-1$

### v2

@@ 4.2. Step 1: the key combinatorial lemma
The following lemma follows from a quite general argument described in a parallel paper, see  [4, Remark 5.4].
@@
Proof.
Instead of giving a proof of this lemma, we rather make a precise connection to the notation and statements in [4] thus expanding [4, Remark 5.4].
The class $\Psi(v)$ in our setting, which appears in the definitions of $B_{g,n}^{m}$ and $\Xi_{g,n}^{m}$, is replaced in [4] by a different class denoted $\mathbb{\Pi}(v)$. We show that the replacement
$\Psi(v)\leftrightarrow\mathbb{\Pi}(v)$
induces precisely to the replacements
[…]
and
(4.4) $\Xi_{g,n}^{m}$ $\leftrightarrow{\Upsilon\Upsilon}_{g,n}^{m}$
in [4].
[…]
The replacement (4.4) requires some further explanation. Indeed, the class $\Xi^{m}_{g,n}$ is defined in Equation (3.6) as a sum over pre-stable star rooted trees, while the class ${\Upsilon\Upsilon}^{m}_{g,n}$ is defined in [4, Equation (4.3)] as the sum over stable star rooted trees with at least two vertices and two extra summands, $\delta_{m,1}{\mathbb{D}}_{g,n+1}$ and $\mathbb{\Pi}^{m}_{g,n}$. In fact, for $2g-2+m+n$ there are three possible cases of pre-stable star rooted trees:
  1. (1)
    The root is stable and all other vertices are unstable. Under the push-forward in the definition of $\Xi^{m}_{g,n}$ this gives the class $\Psi(v_{r})$ for the stable rooted tree that consists of just one vertex $v_{r}$. This matches the exceptional summand $\mathbb{\Pi}^{m}_{g,n}$ in the notation of [4] under the replacement $\Psi(v)\leftrightarrow\mathbb{\Pi}(v)$.
  2. (2)
    The root is not stable. In this case $m=1$ and the star structure implies that there is just one more vertex in the corresponding pre-stable star rooted tree, decorated by ${\mathbb{D}}_{g,n+1}$. This exactly matches the other exceptional summand in [4, Equation (4.3)].
  3. (3)
    The root is stable and there is at least one more stable vertex. In this case we can stabilize the pre-stable star rooted tree forgetting the unstable vertices, and we obtain a one-to-one correspondence with the stable rooted trees of level exactly $1$ in the notation of [4, Equation (4.3)].
[…]
$\left(B_{g,n}^{m}-\delta_{m,1}A_{g,n}^{1}\right)_{d}-\left(-1\right)^{d}\left(\Xi_{g,n}^{m}\right)_{d},$
where $\left(B_{g,n}^{m}-\delta_{m,1}A_{g,n}^{1}\right)_{d}$ corresponds to the analogue of the first term on the right-hand side of [4, Equation (5.3)], and $\left(-1\right)^{d}\left(\Xi_{g,n}^{m}\right)_{d}$ corresponds to the trivial graph contribution from the final term in [4, Equation (5.3)]. Since the full structure of this combinatorial identity is rather cumbersome and irrelevant to the proof in the present paper, we just summarize its key properties that are stated in the lemma. ∎
@@
As we see from the statement of Lemma 4.1, is it more convenient to use the class $\tilde{\Xi}^{m}_{g,n}\coloneqq\sum_{d\geq 0}(-1)^{d}(\Xi^{m}_{g,n})_{d}$.
@@ 4.3. Step 2: equivalence between the master relation and DR/DZ
It follows by induction using the key lemma that the master relation (Conjecture 3.4) is equivalent to the generalized $A=B$ relations (Conjecture 2.6), more precisely:
@@ 4.4. Step 3: proof of the master relation in the Gorenstein quotient
We prove Theorem 3.5 for $\alpha=\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}$ which implies Theorem 3.5 for any $\alpha$ by Lemma 3.6.
@@
    (4.6) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+1}}\psi_{n+1}^{b}\left(B^{1}_{g^{\prime},n^{\prime}}-A^{1}_{g^{\prime},n^{\prime}}\right)\leq 2g^{\prime}-1$